### v1

@@ Abstract
Urban mobility systems are transitioning toward electric, on-demand services, creating operational challenges for fleet management under energy and service-quality constraints. The Electric Dial-a-Ride Problem (E-DARP) extends the classical dial-a-ride problem by incorporating limited battery capacity and nonlinear charging dynamics, increasing computational complexity and limiting the scalability of exact methods for real-time use. This paper proposes a deep reinforcement learning approach based on a graph neural network encoder and an attention-driven route construction policy. By operating directly on edge attributes such as travel time and energy consumption, the method captures non-Euclidean, asymmetric, and energy-dependent routing costs in real road networks. The learned policy jointly optimizes routing, charging, and service quality without relying on Euclidean assumptions or handcrafted heuristics. The approach is evaluated on two case studies using ride-sharing data from San Francisco. On benchmark instances, the method achieves solutions within 0.4% of best-known results while reducing computation times by orders of magnitude. A second case study considers large-scale instances with up to 250 request pairs, realistic energy models, and nonlinear charging. On these instances, the learned policy outperforms Adaptive Large Neighborhood Search (ALNS) by 9.5% in solution quality while achieving 100% service completion, with sub-second inference times compared to hours for the metaheuristic. Finally, sensitivity analyses quantify the impact of battery capacity, fleet size, ride-sharing capacity, and reward weights, while robustness experiments show that deterministically trained policies generalize effectively under stochastic conditions.
@@ 1 Introduction
The Electric Dial-a-Ride Problem (E-DARP) extends the classical Dial-a-Ride Problem by incorporating electric vehicle constraints such as limited driving range and charging requirements. Originally introduced as a generalization for electric and autonomous fleets [BONGIOVANNI2019436], the E-DARP is characterized by its focus on passenger transportation with user-centric service quality constraints: time windows that bound when service can begin, maximum ride time limits that constrain how long passengers spend in vehicles, and precedence requirements ensuring pickups occur before deliveries. These constraints, combined with vehicle capacity limits and battery feasibility requirements, create a computationally challenging problem. Even without electric constraints, the classical DARP is NP-hard, and a single vehicle serving $n$ requests admits $(2n)!/2^{n}$ distinct precedence-feasible orderings—approximately $2.38\times 10^{18}$ sequences for just $n=10$ requests [cordeau2003tabu]. The integration of nonlinear charging dynamics and energy-dependent routing further compounds this complexity.
@@
Exact optimization methods for dial-a-ride problems have progressed significantly, with branch-and-cut [cordeau2006branch, braekers2014exact] and branch-and-price approaches [ropke2009branch, gschwind2015effective] solving classical DARP instances with up to 8 vehicles and 96 requests. Recent work has extended these methods to electric vehicle settings: [goeke2019granular] developed granular tabu search for the E-PDPTW with partial recharging, [liu2023robust] proposed robust optimization for demand uncertainty, and most recently, [su2024branch] introduced the first branch-and-price algorithm specifically for E-DARP using fragment-based representations and novel labeling techniques. Despite these advances, the combinatorial complexity introduced by time windows, capacity limits, ride-time constraints, and nonlinear charging dynamics renders exact methods computationally prohibitive for large-scale, real-time applications requiring sub-second response times. These scalability challenges have motivated growing interest in learning-based approaches. Deep reinforcement learning (DRL) has emerged as a promising paradigm for combinatorial optimization, offering an unsupervised learning approach that discovers effective policies through trial-and-error interaction rather than requiring labeled solutions from existing solvers [bengio2021machine]. Foundational work on attention mechanisms [vinyals2015pointer, bello2017neural] demonstrated viability for vehicle routing [nazari2018reinforcement, kool2019attention]. The POMO framework [kwon2020pomo] showed that learned policies can match or exceed classical heuristics by exploiting multiple solution trajectories during training—an approach particularly relevant to E-DARP where sequential decision structure and robust policy learning align well with POMO’s design. Extensions to pickup-and-delivery problems have introduced heterogeneous attention mechanisms respecting precedence constraints [li2021heterogeneous] and symmetric architectures exploiting paired location structure [luo2024solving]. However, these methods primarily address basic PDP or PDPTW without energy constraints, leaving the integration of battery management and charging decisions largely unexplored in the learning-based literature.
Despite advances in both exact methods and DRL, the application of learning-based approaches to energy-aware dial-a-ride problems remains limited. Supervised learning approaches require large datasets of high-quality solutions generated by existing solvers, creating a computational bottleneck and limiting applicability to problem variants where efficient solvers exist [vinyals2015pointer]. Most DRL research addresses either conventional DARP or electric vehicle routing without ride-time constraints [schneider2014electric, keskin2016partial]. Notable exceptions include [basso2022dynamic], who apply safe reinforcement learning to dynamic stochastic electric vehicle routing, though without the passenger-centric constraints of dial-a-ride. Meanwhile, optimization-based E-DARP literature [BONGIOVANNI2019436, su2024branch] provides valuable benchmarks but requires computation times unsuitable for real-time deployment and does not scale.
@@
At the same time, applying DRL to E-DARP presents distinct challenges. First, the combination of passenger time windows, vehicle capacity, ride-time limits, and battery constraints creates a complex constraint structure that standard neural architectures struggle to represent effectively. Second, and crucially, the E-DARP with energy-dependent routing is inherently a non-Euclidean problem: energy consumption depends on factors such as vehicle load, terrain, and traffic conditions that cannot be captured by coordinate-based distance calculations. Standard attention-based architectures for vehicle routing [kool2019attention, kwon2020pomo] typically embed node locations as Euclidean coordinates and derive edge features (distances, travel times) from these coordinates. This approach fundamentally cannot represent asymmetric travel times, asymmetric real world energy consumption, or real-world road network structures which are asymmetric. Consequently, node-based graph neural networks and standard attention mechanisms are ill-suited for learning effective policies in energy-aware routing problems. Third, while DRL has shown promise on synthetic Euclidean benchmarks, scalability to operationally meaningful instance sizes with realistic, non-Euclidean cost structures remains underexplored. Most existing work evaluates on instances with fewer than 100 requests, leaving a gap between academic benchmarks and the larger problems encountered in practice. Finally, the interpretability and transparency of learned policies presents an ongoing challenge. Neural network-based approaches are often criticized as black-box models, making it difficult to understand why specific routing or charging decisions are made [rudin2019stop, molnar2020interpretable]. In safety-critical transportation applications, this opacity can hinder trust, regulatory acceptance, and debugging of failure cases [arrieta2020explainable]. While attention mechanisms offer some degree of interpretability by revealing which problem elements influence decisions [vaswani2017attention, kool2019attention], systematic frameworks for explaining DRL policies in vehicle routing remain underdeveloped [glanois2024survey]. Recent work has made progress through post-hoc explanation methods [puiutta2020explainable], mechanistic interpretability of neural routing solvers [narad2025mechanistic], and interpretable policy architectures [silva2020optimization], suggesting that learning-based methods are becoming increasingly amenable to interpretation, though this remains an active area of research beyond the scope of the present work.
@@
This work applies the GREAT (Graph Edge Attention Network) encoder architecture [lischka2025greatarchitectureedgebasedgraph] to the E-DARP within an unsupervised deep reinforcement learning framework. Unlike supervised approaches that require optimal or near-optimal solutions for training, our method learns routing policies entirely through reward-driven exploration, eliminating dependence on existing solvers and enabling application to problem scales and variants where generating training labels would be computationally prohibitive. GREAT is specifically designed to address the non-Euclidean nature of real-world routing problems. Whereas conventional node-based architectures derive edge information from node coordinates, GREAT operates directly on edges: information is compared between and passed along edges sharing endpoints, with edge features serving as the primary inputs to the attention mechanism. This edge-centric design is critical for energy-dependent routing, where edge features such as travel times and energy consumption are given directly from road network data rather than being computed from Euclidean coordinates. The architecture can naturally represent asymmetric costs (different energy consumption in each direction), and asymmetric graph structures. For the E-DARP, routing decisions correspond to selecting edges (transitions between pickup locations, delivery locations, and charging stations), making GREAT’s edge-based formulation a natural fit. The representation captures the problem’s intrinsic structure, encoding battery constraints through energy-consumption edge features, time windows through temporal feasibility features, and ride-time restrictions through service-time attributes, without requiring auxiliary constraint-encoding mechanisms or coordinate-based approximations.
@@
    Unsupervised learning framework with edge-based neural architecture. We demonstrate that the GREAT encoder architecture, combined with deep reinforcement learning, enables fully unsupervised learning of E-DARP routing policies. The model discovers effective strategies through reward-driven exploration without requiring labeled solutions from optimization algorithms. By processing routing problems through edge-level attention rather than node-based representations and operating directly on edge features (travel times, distances, energy consumption), the architecture naturally handles asymmetric costs and non-Euclidean network structures that characterize real-world electric vehicle routing. This design choice, combined with an attention-based decoder and POMO training, enables the policy to learn sophisticated routing and charging strategies.
@@
    Comprehensive sensitivity analysis across operational, algorithmic, and environmental factors. We conduct systematic analyses examining: (i) operational attributes, including battery capacity (20 to 40 kWh), fleet size (4 to 8 vehicles), and ride-sharing capacity (1 to 3 passengers), revealing that larger batteries achieve 6 to 15% higher profits, 8-vehicle fleets are necessary for guaranteed coverage, and ride-sharing introduces predictable lateness trade-offs; (ii) reward function weights, demonstrating that adequate completion incentives are essential for full service coverage while excessive operational cost penalties lead to conservative behavior and incomplete service; (iii) hyperparameters, identifying that smaller neural architectures (hidden dimension 64 to 128) match or outperform larger models while converging faster; and (iv) robustness under stochastic travel times and energy consumption, showing that deterministic training produces policies generalizing effectively to 10% operational uncertainty.
Together, these contributions demonstrate the potential of learning-based, edge-centric models to tackle large-scale, energy-aware routing problems efficiently and adaptively, providing a practical framework for real-world autonomous fleet deployment. The remainder of this paper is organized as follows. We begin by formulating the E-DARP as a Markov Decision Process and providing a detailed mathematical problem statement including constraints and objectives (Section 2). The methodology is then presented, describing the GREAT encoder architecture, the attention-based decoder, feasibility masking, and the reward function design (Section 3). Next, we detail the experimental setup, including problem instance generation, training procedures, and baseline methods for both benchmark and large-scale case studies (Section 4). Comprehensive results and analysis follow, covering benchmark performance comparison against exact methods, large-scale curriculum learning evaluation, sensitivity analysis across battery sizes, fleet configurations, and ride-sharing capacities, hyperparameter analysis, reward weight sensitivity, and robustness evaluation under stochastic conditions (Section 5). Finally, we conclude with a summary of findings, discussion of limitations, and directions for future research (Section 6).
@@ 3.1 Problem Encoder
The GREAT encoder consist of several stacked individual GREAT layers (in particular, we use the GREAT NB layer version of [lischka2025greatarchitectureedgebasedgraph]). Each layer consists of an edge-based multi-head attention sublayer, a feed forward layer, residual layers and normalizations. We refer the reader to [lischka2025greatarchitectureedgebasedgraph] for details. Here it suffices to say that at each layer $\ell$, the representation of edge(i,j)(i,j) is updated by attending to neighboring edges that share an endpoint with it. Specifically, we compute
@@ 4 Experimental Setup
This section evaluates the proposed approach on two case studies based on real-world ride-sharing data from San Francisco. Both case studies use transportation requests derived from the Uber GPS dataset [uber2011gps], which contains approximately 1.2 million GPS logs recorded every 4 seconds from active Uber vehicles during one week in San Francisco in 2011. After removing invalid records and extracting pickup and dropoff locations, the processed dataset contains about 25,000 trips with daily request volumes ranging from 2,000 to 6,000. The dataset is split into training, validation, and test sets.
4.1 Problem Instances and Training
@@
The benchmark objective differs from the formulation in Section 2.3. Rather than minimizing energy, waiting time, and tardiness, the benchmark minimizes travel time and tardiness with weights $w_{1}=0.75$ for travel time and $w_{3}=0.25$ for tardiness ($w_{2}=0$ for waiting time). Additionally, the benchmark formulation does not include the completion bonus term, implicitly assuming all requests must be served. We adopt these objective weights and constraint specifications to ensure direct comparability with published results.
Training instances for Case Study 1 are generated following these same specifications, sampling requests from the processed Uber GPS dataset. The neural network is trained using REINFORCE [williams1992simple] with POMO augmentation [kwon2020pomo] to increase variance and improve solution quality. To evaluate generalization across problem scales, we employ mixed-size training with instances containing $n\in\{16,24,32,40,50\}$ request pairs, sampled uniformly during training. The training dataset consist of 5000 graphs, the validation dataset of 1000 graphs, and the test dataset of the benchmark cases.

### v2

@@ Abstract
Urban mobility systems are transitioning toward electric, on-demand services, creating operational challenges for fleet management under energy and service-quality constraints. The Electric Dial-a-Ride Problem (E-DARP) extends the classical dial-a-ride problem by incorporating limited battery capacity and nonlinear charging dynamics, increasing computational complexity and limiting the scalability of exact methods for real-time use. This paper proposes a deep reinforcement learning approach based on an edge-centric graph neural network encoder and an attention-driven route construction policy. By operating directly on edge attributes such as travel time and energy consumption, the method captures non-Euclidean, asymmetric, and energy-dependent routing costs in real road networks. The learned policy jointly optimizes routing, charging, and service quality without relying on Euclidean assumptions or handcrafted heuristics. The approach is evaluated on two case studies using ride-sharing data from San Francisco. On benchmark instances, the method achieves solutions within 0.4% of best-known results while reducing computation times by orders of magnitude. A second case study considers large-scale instances with up to 250 request pairs, realistic energy models, and nonlinear charging. On these instances, the learned policy outperforms Adaptive Large Neighborhood Search (ALNS) by 9.5% in solution quality while achieving 100% service completion, with inference times under 10 seconds compared to hours for the metaheuristic. Finally, sensitivity analyses quantify the impact of battery capacity, fleet size, ride-sharing capacity, and reward weights, while robustness experiments show that deterministically trained policies generalize effectively under stochastic conditions.
@@ 1 Introduction
The Electric Dial-a-Ride Problem (E-DARP), introduced by [BONGIOVANNI2019436], formalizes these challenges by extending the classical Dial-a-Ride Problem with electric vehicle constraints such as limited driving range and charging requirements. In this problem, a fleet of electric vehicles operates from one or more depots to serve passenger transportation requests, where each request consists of picking up passengers at an origin and delivering them to a destination. The problem incorporates user-centric service quality constraints: time windows bound when service can begin and end, maximum ride time limits constrain how long passengers spend in vehicles, and ordering constraints ensure pickups occur before deliveries. Additionally, vehicles have limited battery capacity and must visit charging stations when needed. These constraints, combined with vehicle capacity limits, create a computationally challenging problem. Even without electric constraints, the classical DARP is NP-hard, and a single vehicle serving $n$ requests admits $(2n)!/2^{n}$ possible solutions, approximately $2.38\times 10^{18}$ for just $n=10$ requests [cordeau2003tabu].The integration of nonlinear charging dynamics and energy-dependent routing further compounds this complexity.
@@
Despite this complexity, exact optimization methods for dial-a-ride problems have made significant progress. Branch-and-cut [cordeau2006branch, braekers2014exact] and branch-and-price approaches [ropke2009branch, gschwind2015effective] can now solve classical DARP instances with up to 8 vehicles and 96 requests. Recent work has extended these methods to electric vehicle settings, including branch-and-cut [BONGIOVANNI2019436], granular tabu search [goeke2019granular], robust optimization for demand uncertainty [liu2023robust], and branch-and-price with fragment-based representations [su2024branch]. Despite these advances, the combinatorial complexity introduced by time windows, capacity limits, ride-time constraints, and nonlinear charging dynamics renders exact methods computationally prohibitive for large-scale, real-time applications requiring sub-second response times. These scalability challenges have motivated growing interest in learning-based approaches. Deep reinforcement learning (DRL) has emerged as a promising paradigm for combinatorial optimization, offering an unsupervised learning approach that discovers effective policies through trial-and-error interaction rather than requiring labeled solutions from existing solvers [bengio2021machine]. Foundational work on attention mechanisms [vinyals2015pointer, bello2017neural] demonstrated effectiveness for vehicle routing [nazari2018reinforcement, kool2019attention]. The POMO framework [kwon2020pomo] showed that learned policies can match or exceed classical heuristics by exploiting multiple solution trajectories during training, an approach particularly relevant to E-DARP where sequential decision structure and robust policy learning align well with POMO’s design. Extensions to pickup-and-delivery problems have introduced heterogeneous attention mechanisms respecting ordering constraints [li2021heterogeneous] and symmetric architectures exploiting paired location structure [luo2024solving]. However, these methods primarily address basic PDP or PDPTW without energy constraints, leaving the integration of battery management and charging decisions largely unexplored in the learning-based literature.
@@
Despite advances in both exact methods and DRL, learning-based approaches have not been applied to the full E-DARP. Supervised learning requires high-quality solutions from existing solvers, limiting applicability to problems where such solvers exist [vinyals2015pointer]. Most DRL research addresses either conventional DARP without energy constraints or electric vehicle routing without passenger-centric requirements such as ride-time limits [schneider2014electric, keskin2016partial, basso2022dynamic]. This leaves a gap: no learning-based method currently handles the combination of battery management, charging decisions, and user-centric service constraints that defines the E-DARP. At the same time, applying DRL to E-DARP presents distinct challenges. First, the combination of passenger time windows, vehicle capacity, ride-time limits, and battery constraints creates a complex constraint structure that standard neural architectures struggle to represent effectively. Second, and crucially, the E-DARP with energy-dependent routing is inherently non-Euclidean and asymmetric: traveling from location A to B differs from B to A due to factors such as elevation changes, one-way streets, and traffic patterns. These direction-dependent costs cannot be captured by coordinate-based distance calculations. Standard attention-based architectures for vehicle routing [kool2019attention, kwon2020pomo] typically embed node locations as Euclidean coordinates and derive edge features (distances, travel times) from these coordinates. This approach fundamentally cannot represent asymmetric travel times, asymmetric real world energy consumption, or real-world road network structures which are asymmetric. Consequently, node-based graph neural networks and standard attention mechanisms are ill-suited for learning effective policies in energy-aware routing problems. Third, while DRL has shown promise on synthetic Euclidean benchmarks, scalability to operationally meaningful instance sizes with realistic, non-Euclidean cost structures remains underexplored. Most existing work evaluates on instances with fewer than 100 requests, leaving a gap between academic benchmarks and the larger problems encountered in practice. Finally, the interpretability and transparency of learned policies presents an ongoing challenge. Neural network-based approaches are often criticized as black-box models, making it difficult to understand why specific routing or charging decisions are made [rudin2019stop, molnar2020interpretable]. In safety-critical transportation applications, this opacity can hinder trust, regulatory acceptance, and debugging of failure cases [arrieta2020explainable]. While attention mechanisms offer some degree of interpretability by revealing which problem elements influence decisions [vaswani2017attention, kool2019attention], systematic frameworks for explaining DRL policies in vehicle routing remain underdeveloped [glanois2024survey]. Recent work has made progress through post-hoc explanation methods [puiutta2020explainable], mechanistic interpretability of neural routing solvers [narad2025mechanistic], and interpretable policy architectures [silva2020optimization], suggesting that learning-based methods are becoming increasingly amenable to interpretation, though this remains an active area of research beyond the scope of the present work.
@@
This work applies the GREAT (Graph Edge Attention Network) encoder architecture [lischka2025greatarchitectureedgebasedgraph] to the E-DARP within an unsupervised deep reinforcement learning framework. Unlike supervised approaches that require optimal or near-optimal solutions for training, our method learns routing policies entirely through reward-driven exploration, eliminating dependence on existing solvers. GREAT is specifically designed for non-Euclidean routing problems. Conventional node-based architectures derive edge information from node coordinates, but GREAT operates directly on edges: edge features serve as the primary inputs to the attention mechanism, and information propagates between edges that share endpoints. This design naturally represents asymmetric costs and direction-dependent energy consumption. For E-DARP, routing decisions correspond to selecting edges between pickup locations, delivery locations, and charging stations, making the edge-based formulation a natural fit.
@@
    Unsupervised learning framework with edge-based graph neural architecture. We demonstrate that the GREAT encoder architecture, combined with deep reinforcement learning, enables fully unsupervised learning of E-DARP routing policies. The model discovers effective strategies through reward-driven exploration without requiring labeled solutions from optimization algorithms. By processing routing problems through edge-level attention rather than node-based representations and operating directly on edge features (travel times, distances, energy consumption), the architecture naturally handles asymmetric costs and non-Euclidean network structures that characterize real-world electric vehicle routing. This design choice, combined with an attention-based decoder and POMO training, enables the policy to learn sophisticated routing and charging strategies.
@@
    Comprehensive sensitivity analysis. We conduct systematic analyses of operational factors (battery capacity, fleet size, ride-sharing capacity), reward function weights, neural architecture hyperparameters, and robustness under stochastic conditions. Key findings include that larger batteries improve profits by 6 to 15%, adequate completion incentives are essential for full service coverage, smaller architectures match or outperform larger models, and deterministically trained policies generalize effectively to 10% operational uncertainty.
Together, these contributions demonstrate that learning-based, edge-centric models can tackle large-scale, energy-aware routing problems efficiently and adaptively, providing a practical framework for real-world autonomous fleet deployment. The remainder of this paper is organized as follows. We begin with an introductory example illustrating the problem’s computational challenges, then formulate the E-DARP as a Markov Decision Process and provide a detailed mathematical problem statement including constraints and objectives (Section 2). The methodology is then presented, describing the GREAT encoder architecture, the attention-based decoder, feasibility masking, and the reward function design (Section 3). Next, we detail the experimental setup, including problem instance generation, training procedures, and baseline methods for both benchmark and large-scale case studies (Section 4). Comprehensive results and analysis follow, covering benchmark performance comparison against exact methods, large-scale curriculum learning evaluation, sensitivity analysis across battery sizes, fleet configurations, and ride-sharing capacities, hyperparameter analysis, reward weight sensitivity, and robustness evaluation under stochastic conditions (Section 5). Finally, we conclude with a summary of findings, discussion of limitations, and directions for future research (Section 6).
@@ 3.1 Problem Encoder
The GREAT encoder consists of several stacked individual GREAT layers (in particular, we use the GREAT NB layer version of [lischka2025greatarchitectureedgebasedgraph]). Each layer consists of an edge-based multi-head attention sublayer, a feed forward layer, residual layers and normalizations. We refer the reader to [lischka2025greatarchitectureedgebasedgraph] for details. Here it suffices to say that at each layer $\ell$, the representation of edge(i,j)(i,j) is updated by attending to neighboring edges that share an endpoint with it. Specifically, we compute
@@ 4 Experimental Setup
This section evaluates the proposed approach through two complementary case studies based on real-world ride-sharing data from San Francisco. Case Study 1 uses established benchmark instances from the E-DARP literature [BONGIOVANNI2019436] with 16–50 requests in the Civic Center district, enabling direct comparison with exact optimization methods and published results. Case Study 2 addresses larger instances with 100–250 requests across a broader San Francisco service area, incorporating realistic energy consumption modeling and more realistic electric vehicle specifications to demonstrate scalability beyond the reach of exact methods.
Both case studies use transportation requests derived from the Uber GPS dataset [uber2011gps], which contains approximately 1.2 million GPS logs recorded every 4 seconds from active Uber vehicles during one week in San Francisco in 2011. After removing invalid records and extracting pickup and dropoff locations, the processed dataset contains about 25,000 trips with daily request volumes ranging from 2,000 to 6,000.
4.1 Case Study 1: Benchmark Instances
@@
The benchmark objective differs from the formulation in Section 2.3. Rather than minimizing energy, waiting time, and lateness, the benchmark minimizes travel time and excess travel time with weights $w_{1}=0.75$ for travel time and $w_{2}=0.25$ for excess travel time. Additionally, the benchmark formulation does not include the completion bonus term, implicitly assuming all requests must be served. We adopt these objective weights and constraint specifications to ensure direct comparability with published results.
A key difference between the benchmark formulation and our learned approach concerns the treatment of charging decisions. In the original benchmark, charging time at each station visit is a continuous decision variable, allowing the optimization to determine the exact amount of energy to charge based on future route requirements [BONGIOVANNI2019436]. Our reinforcement learning framework, however, requires a fixed charging time or charging threshold and cannot directly optimize continuous charging durations within the policy network. To address this, we fix the charging time to a predetermined duration at each charging station visit. Setting a relatively small fixed charging time provides finer-grained control to the learned policy: the agent can effectively modulate total charging by deciding how many times to visit charging stations, including consecutive visits to the same station, rather than how long to charge per visit. This discretization trades some solution flexibility for compatibility with our action-based learning framework while still enabling effective battery management strategies. However, this approach may yield suboptimal solutions in scenarios where charging is required, as the fixed duration cannot adapt to the precise energy needs of each route, potentially resulting in either insufficient charge or unnecessary time spent at stations.
The two case studies use different objective functions to address distinct evaluation goals. Case Study 1 adopts the benchmark objective to enable direct comparison with published exact methods. This isolates the question of whether learned policies can match optimization approaches on established instances. Case Study 2 uses the full E-DARP objective from Section 3.4, which better reflects operational priorities for electric fleets such as energy efficiency and service completion under battery constraints. Although this prevents direct numerical comparison between case studies, both objectives share the core structure of minimizing travel costs subject to time window and capacity constraints. The strong performance across both formulations suggests the GREAT architecture captures the underlying routing structure rather than being specific to one objective.
Training instances for Case Study 1 are generated following these same specifications, sampling requests from the processed Uber GPS dataset. The neural network is trained using REINFORCE [williams1992simple] with POMO [kwon2020pomo] to increase variance and improve solution quality. To evaluate generalization across problem scales, we employ mixed-size training with instances containing $n\in\{16,24,32,40,50\}$ request pairs, sampled uniformly during training. The training dataset consists of 5000 graphs, the validation dataset of 1000 graphs, and the test dataset of the benchmark cases.
4.2 Case Study 2: Large-Scale Realistic Instances